OpenStudy (anonymous):

f(x)= 2x+3 g(x) = 5 - 4x x=4 Solve for f(g(x)) and g(f(x))

zepdrix (zepdrix):

\[\Large f(x)=2x+3 \qquad \qquad g(x)= 5-4x\qquad\qquad x=4\] Hmm I'm trying to figure out why they included x=4. They probably meant that they want us to solve for, \(\large f(g(4)) \qquad\text{and}\qquad g(f(4))\)

zepdrix (zepdrix):

\[\Large f(\color{royalblue}{x})=2\color{royalblue}{x}+3\]For the first one, we want to plug g(x) in for our blue x that's in f.\[\Large f(\color{royalblue}{g(x)})=2\color{royalblue}{g(x)}+3\]Which gives us,\[\Large f(\color{royalblue}{g(x)})=2\color{royalblue}{(5-4x)}+3\]

zepdrix (zepdrix):

If we wanted to evaluate this at x=4, we have,\[\Large f(\color{royalblue}{g(4)})=2\color{royalblue}{(5-4\cdot4)}+3\]

zepdrix (zepdrix):

What do you think Crys? :x too confusing?

OpenStudy (anonymous):

No I get the concept! :)

OpenStudy (anonymous):

Will the answer be 11? @zepdrix

zepdrix (zepdrix):

For f(g(4))? Hmm it looks like it's -19. \[\Large f(4)=2\cdot4+3=11\] Did you calculate that out by mistake? :o

OpenStudy (anonymous):

yea i used calculator

zepdrix (zepdrix):

\[\Large \color{royalblue}{g(4)=5-4\cdot4=-11}\]\[\Large \color{royalblue}{g(4)=-11}\] We want to plug this value in for f(x).

zepdrix (zepdrix):

\[\Large f(\color{royalblue}{g(4)}) \qquad=\qquad f(\color{royalblue}{-11})\]

zepdrix (zepdrix):

\[\Large f(-11)=2(-11)+3\]

OpenStudy (anonymous):

Ok I see how I get -19 now!

zepdrix (zepdrix):

Function notation so confusing! :) Takes a little getting used to. Try to find your g(f(4)) also

OpenStudy (anonymous):

I just did pemdas

OpenStudy (anonymous):

Ok Ill let u know what I get

OpenStudy (anonymous):

So will it start out with 5-4x(2x+3)

zepdrix (zepdrix):

5-4(2x+3) careful, i see an extra x in there! :O we `replace` the x with that big mess of parentheses.

OpenStudy (anonymous):

how is that?

OpenStudy (anonymous):

ohhhh i seee lol

OpenStudy (anonymous):

distrubuted and got 10x +15 -8x-12

OpenStudy (anonymous):

do i sub the 4 now?

zepdrix (zepdrix):

oh sorry i disappeared there +_+ Hmmm i don't understand how you got that. Let's see,\[\Large f(x)=5-4x\]\[\Large f(g(x))=5-4(2x+3)\]

zepdrix (zepdrix):

Distributing the -4 gives us,\[\Large f(g(x))=5-8x-12\]

OpenStudy (anonymous):

then u take the 5- 12?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

ya -8x-7